OpenStudy (anonymous):

Evaluate the erxpression for a=6 and b=14 1. 2a 3b 2. 2/a * 3/b Help please?!

OpenStudy (anonymous):

How about this one? @BangkokGarrett

OpenStudy (anonymous):

is 1. 2a * 3b?????????

OpenStudy (anonymous):

1

OpenStudy (anonymous):

One is 2a + 3b @Krissy1997

OpenStudy (anonymous):

(2*6)+(3*14)= (2/6)*(3/14)=

OpenStudy (anonymous):

1. 54 2. 07142857142

OpenStudy (anonymous):

GTG